The area of a rug is 48 square feet and the width is 2 feet less than the length, find the width of the carpet. I know the answer is 6 feet. BUT HOW DO YOU GET 6???
Answer by ewatrrr(24785)

 
Hi,
The area of a rug is 48 square feet and the width is 2 feet less than the length, find the width of the carpet.

Let x represent the length and (x-2) represent the width
Question states*****
x(x-2) = 48 ft^2
Solving for x
x^2 - 2x - 48 = 0
(x+6)(x-8) = 0
(x+6) = 0, x = -6 (throwing out negative solution for unit measure)
(x-8) = 0 , x = 8', the length and the width is 6' (x-2)
CHECKING our Answer***
( 8')(6') = 48 ft^2